11. How many moles, how many molecules, and how many stoms are present in 50.0 of a) F; b) H0 c) CCL d) BCI 033 6 38,0 CCM
Answer #1


Number of moles of substance = MASS / MOLAR MASS

1 mole of any substance any contains 6.023*10^23 substances


1. Number of moles of F2 = 50/38 = 1.32 moles.

Number of molecules of F2 present in 1.32 moles = 1.32*6.023*10^23 molecules

                                                = 7.95*10^23 molecules

Number of atoms of F present in 7.95*10^23 F2 molecules = 2* 7.95*10^23 = 1.59*10^24 atoms.

2. Number of moles of H2O = 50/18 = 2.78 moles.

Number of molecules of H2O present in 2.78 moles = 2.78*6.023*10^23 molecules

                                                = 1.68*10^24 molecules

Number of atoms in 1.68*10^24 H2O molecules = 3*1.68*10^24 = 5*10^24 atoms.

3. Number of moles of CCl4 = 50/154 = 0.32 moles.

Number of molecules of CCl4 present in 0.32 moles = 0.32*6.023*10^23 molecules

                                                = 1.92*10^23 molecules

Number of atoms in 1.92*10^23 CCl4 molecules = 5*1.92*10^23   = 9.6*10^23 atoms.

4. Number of moles of BCl3 = 50/117 = 0.43 moles.

Number of molecules of BCl3 present in 0.43 moles = 0.43*6.023*10^23 molecules

                                                = 2.59*10^23 molecules

Number of atoms in 2.59*10^23 BCl3 molecules = 4*2.59*10^23   = 1.03*10^24 atoms